formula for equivalent weight of acids
 

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
The weight of a compound that contains ONE EQUIVALENT of a proton (for acid)

Examples:
(1) H2SO4 + 2OH-= 2H2O + SO42-
Molecular weight of sulfuric acid (H2SO4) = 98.07 g/mol
Theoretical Equivalent weight of sulfuric acid (H2SO4) = 49.03 g/equivalent of H+
Reasoning:
98.07 g/mol H2SO4 * (1 mole H2SO4 / 2 equivalents of H+) = 49.03 g/equivalent of H+
In short
Equivalent weight,E=\frac{molecular weight}{basicity}